# references.bib

### v3

@@ Appendix A Mathematical comparison of PPI and post-stratified area estimator
We show that when the number of map product points $N$ is large, the PPI and post-stratified area estimator confidence intervals are similar. In particular, we show (1) that they have approximately the same length when the number of samples is large and (2) they are centered on the same point. These two facts together provide a mathematical explanation for why PPI and the post-stratified estimator approximately agree in all of our experiments.
@@
We now turn to the PPI standard error. We have
$SE_{\text{PPI}}$ $\rightarrow\sqrt{\frac{1}{n}\mathrm{Var}(Y-\lambda\hat{Y})}$ since $N\gg n$
@@
$SE_{\text{post}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
@@
Appendix B Binarizing the NASA Global Forest Cover Change map product for deforestation
In the Brazilian Amazon deforestation use cases, the ground truth points from [bullock2020satellite] have binary deforestation values ($Y=1$ if deforested and $Y=0$ if not deforested between 2000-2015), while the NASA GFCC map product gives canopy cover percentages over time at each pixel.
In order to binarize the continuous map product data, we classify a map pixel as “deforested” if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$. We choose the $-25\%$ threshold by visually comparing the map-derived canopy cover change distributions for deforested and non-deforested ground truth points (Figure 14). Most of the deforested ground truth points have a map canopy cover decrease of more than $25\%$, while most of the non-deforested ground truth points have either a map canopy cover increase, no map canopy cover change, or a map canopy cover decrease of less than $25\%$.

### v4

@@ Appendix F Mathematical comparison of PPI and post-stratified area estimator
We show that when the number of map product points $N$ is large, the PPI and post-stratified area estimator confidence intervals are similar. In particular, we show (1) that they have approximately the same length when the number of sample units is large and (2) they are centered on the same point. These two facts together provide a mathematical explanation for why PPI and the post-stratified estimator approximately agree in all of our experiments.
@@
We now turn to the PPI standard error. We have the standard deviation of the tuned PPI estimator from [angelopoulos2023ppi++]
$\sigma_{\text{PPI}}$ $\rightarrow\sqrt{\frac{1}{n}\mathrm{Var}(Y-\lambda\hat{Y})}$ since $N\gg n$
@@
A standard error estimate $\hat{\sigma}_{\text{PPI}}$ of $\sigma_{\text{PPI}}$ also satisfies the above equations.
@@
$\hat{\sigma}_{\text{post}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
@@
Appendix G Simulating mean reversion bias in maps
In Section 4, we simulate adding noise and nonlinear bias to the map products from Example 4 (forest cover linear regression). In this appendix, we simulate another model of map bias: linear mean reversion. Under this type of bias, the map underestimates high ground truth values and overestimates low ground truth values.
Using the same MOSAIKS dataset as Section 4, we simulate different levels of mean reversion bias in the population map (error-in-$X$), the forest cover map (error-in-$Y$), or both (error-in-both). The elevation dataset is not altered. We compare the resulting 95% coefficient confidence intervals at different levels of simulated map bias using map-only, GT-only, and PPI methods.
Bias
To simulate bias for the ground truth population covariate $X_{\text{pop}}$, we linearly interpolate between $X_{\text{pop}}$ and the mean $\mu_{X_{\text{pop}}}$ of all $X_{\text{pop}}$ values in the full ground truth dataset (67968 points). The resulting simulated map product value at “bias level $c$” is
$X_{\text{pop}}^{\text{map}}=c\mu_{X_{\text{pop}}}+(1-c)X_{\text{pop}}.$
Similarly, for the forest cover response variable $Y$ with mean $\mu_{Y}$, the simulated map product value at each point is
$Y^{\text{map}}=c\mu_{Y}+(1-c)Y.$
We experiment with bias levels $c$ from 0 to 0.8 at increments of 0.1. When $c=0$, the simulated map is the same as the full ground truth map.
Results
The coefficient point estimates and 95% confidence intervals at each bias level are shown in Figure 14. We compare results for the map-only, GT-only, and PPI estimators. (For all experiments in this section, we use $B=200$ bootstrap iterations to construct PPI confidence intervals.)
Similar to Section 4, the GT-only confidence interval stays constant regardless of the bias level, and it is wider than the PPI and map-only intervals due to the small number of ground truth points ($n=500$).
At all bias levels across all three experiments (error in $X_{\text{pop}}$, $Y$, or both), the PPI confidence intervals contain the true coefficients for both covariates. This is expected, as PPI uses the small set of ground truth points to correct for bias in the map product, and the PPI confidence intervals are guaranteed to contain the true coefficient with 95% probability.
At all bias levels, the PPI confidence intervals are significantly smaller than the GT-only intervals. This is because the biased map is a simple linear transformation of the true variable values and thus provides useful information for estimating the regression coefficients. As discussed in Section 4, PPI allows us to conclude that there is a positive association between forest cover and population, while the GT-only interval for $\beta_{\text{pop}}$ contains zero and does not detect this significant association.
In all three experiments, as map bias increases, the map-only estimates become biased for at least one of the coefficients. The map-only confidence intervals are very narrow because the number of map product points is large ($N=67968$), so they fail to contain the true coefficients as the estimates become biased. As mean reversion bias in $X_{\text{pop}}$ increases, the map underestimates high population values and overestimates low population values. As a result, the map-only estimates for $\beta_{\text{pop}}$ increase in magnitude and overestimate the true coefficient, while map-only estimates for $\beta_{\text{elevation}}$ remain unbiased (since we do not add bias to the elevation map). As mean reversion bias in $Y$ (forest cover) increases, the map underestimates high forest cover values and overestimates low forest cover values. As a result, the map-only coefficient estimates for both covariates decrease in magnitude and attenuate toward zero. When there is mean reversion at level $c$ in both forest cover and population, the effects of the two biases cancel out to produce unbiased map-only estimates of $\beta_{\text{pop}}$. However, the map-only estimates of $\beta_{\text{elevation}}$ still decrease in magnitude and attenuate toward zero.
Figure 14: Simulation estimating forest cover linear regression coefficient 95% confidence intervals under different levels of map mean reversion bias. PPI gives unbiased estimates with much narrower confidence intervals than the GT-only method. In each experiment, the map-only coefficient estimate for at least one covariate becomes increasingly biased as map bias increases.